### v2

# EasySteer: A Unified Framework for High-Performance and Extensible LLM Steering

Abstract
Large language model (LLM) steering has emerged as a promising paradigm for controlling model behavior at inference time through targeted manipulation of hidden states, offering a lightweight alternative to expensive retraining. However, existing steering frameworks suffer from critical limitations: computational inefficiency, limited extensibility, and restricted functionality that hinder both research progress and practical deployment. We present EasySteer, a unified framework for high-performance, extensible LLM steering built on vLLM. Our system features modular architecture with pluggable interfaces for both analysis-based and learning-based methods, fine-grained parameter control, pre-computed steering vectors for eight application domains, and an interactive demonstration system. Through deep integration with vLLM’s optimized inference engine, EasySteer achieves 10.8-22.3 speedup over existing frameworks. Extensive experiments demonstrate its effectiveness in overthinking mitigation, hallucination reduction, and other key applications. EasySteer transforms steering from research technique to production-ready capability, establishing critical infrastructure for deployable, controllable language models.
EasySteer: A Unified Framework for High-Performance and Extensible LLM Steering
Haolei Xu1, Xinyu Mei1, Yuchen Yan1, Rui Zhou1, Wenqi Zhang1, Weiming Lu1††thanks: Corresponding author., Yueting Zhuang1, Yongliang Shen111footnotemark: 1 1Zhejiang University {xuhaolei,luwm,syl}@zju.edu.cn Open-source repository: https://github.com/ZJU-REAL/EasySteer Demonstration video: https://www.youtube.com/watch?v=3rRGzZmhrXg
1 Introduction
Large language models (LLMs) have achieved remarkable capabilities, yet controlling their behavior during deployment remains a fundamental challenge (Zhao et al., 2023). Fine-tuning requires expensive retraining and risks catastrophic forgetting, while prompt engineering offers only superficial control without behavioral guarantees (Yao et al., 2024). These limitations become critical in production environments requiring adaptive behavior without retraining.
LLM steering offers a compelling solution through targeted manipulation of hidden states during inference (Turner et al., 2023; Zhang et al., 2026). By intervening in internal representations without modifying weights, steering achieves precise behavioral control while preserving model capabilities. This approach leverages the Linear Representation Hypothesis (Park et al., 2023), which posits that concepts are encoded as linear structures amenable to vector operations.
Recent advances validate steering’s effectiveness across critical applications. Thinking pattern vectors successfully mitigate overthinking in mathematical reasoning (Lin et al., 2025; Liu et al., 2025), preference-based methods achieve personality control (Cao et al., 2024), and simple additive vectors manipulate refusal behaviors (Lee et al., 2024). These successes establish steering as both a practical control mechanism and a tool for mechanistic interpretability.
Despite these advances, practical implementation remains challenging, as steering typically requires modifying the forward propagation process through complex wrappers or hooks, creating significant engineering barriers. Several frameworks have emerged to facilitate steering research, including repeng (Vogel, 2024), pyreft (Wu et al., 2024), and EasyEdit2 (Xu et al., 2025b). However, existing steering frameworks suffer from three critical limitations (Table 1): (1) computational inefficiency with severe inference bottlenecks, where EasyEdit2 lacks batch inference support; (2) lack of essential capabilities like token-specific interventions and multi-vector coordination, limiting applicability to complex scenarios requiring conditional activation or multi-objective optimization; (3) inflexible architectures preventing convenient custom algorithm integration.
| Framework | repeng | pyreft | EasyEdit2 | EasySteer |
| Base Library | Transformers | Transformers | Transformers | vLLM |
| Speedup | 1.0× | 2.1× | N/A | 22.3× |
| Algorithm | Analytical | Learned | Both | Both |
| Granularity | Layer | Layer/Position | Layer/Position | Layer/Position/ Token/Stage |
| Coordination | Multi vector | Single only | Multi vector | Multi vector/ algorithm |
| Extensibility | Limited | Limited | Modular | Modular |
| Multimodal | Text only | Text only | Text & Vision | Text & Vision |
To address these challenges, we present EasySteer, an Apache-2.0 licensed open-source unified framework for high-performance, extensible LLM steering built on vLLM (Kwon et al., 2023). Our system comprises four integrated modules: (1) Steering Vector Generation Module supporting both analysis-based and learning-based methods; (2) Steering Vector Application Module leveraging vLLM’s optimized engine for efficient hidden state intervention with pluggable algorithm interfaces and fine-grained parameter control; (3) Comprehensive Resource Library offering production-ready steering vectors and examples for eight application domains with documented evaluation results; (4) Interactive Demonstration System providing an intuitive web interface for vector extraction, training, inference and chat.
Through deep vLLM integration, EasySteer achieves 10.8-22.3× speedup over existing frameworks while maintaining 81-91% of baseline throughput even under multi-vector configurations. Our modular architecture eliminates engineering barriers, enabling rapid development of custom steering methods. Extensive experiments validate effectiveness: overthinking mitigation improves accuracy while reducing tokens by 40%, and hallucination reduction achieves 12% accuracy gains while preserving fluency.
2 Related Work
Model Control Paradigms.
Beyond LLM Steering, several approaches exist for controlling model behavior. Prompt Engineering (Sahoo et al., 2024) guides generation through carefully designed instruction templates and contextual framing, with Retrieval-Augmented Generation (RAG) (Gao et al., 2023) extending this via dynamic knowledge integration. Fine-tuning methods adapt model behavior through weight updates, ranging from full parameter updates to Parameter-Efficient Fine-Tuning (PEFT) (Han et al., 2024) approaches like LoRA (Hu et al., 2022). Model Editing techniques (Zhang et al., 2024), including ROME (Meng et al., 2022a) and MEMIT (Meng et al., 2022b), enable precise knowledge updates by targeting specific parameters encoding factual information. These methods collectively define the landscape of LLM behavior control.
Mechanistic Interpretability.
Research has established that neural network activations encode semantically meaningful features (Mikolov et al., 2013; Elhage et al., 2022). Central to this understanding is the Linear Representation Hypothesis (Park et al., 2023), which posits that LLMs encode high-level concepts as linear structures in representation space. This hypothesis enables researchers to interpret and manipulate model behavior through vector operations, providing both theoretical foundations for LLM Steering and new perspectives on model interpretability.
3 Formalization of LLM Steering
Consider an L-layer language model processing input sequence . Let denote the hidden state at layer and position , where represents the hidden dimension and denotes the sequence length. We formalize LLM steering as an inference-time transformation function . When a specific condition is satisfied, this function maps to a steered representation . The condition is determined by contextual factors or internal model states and triggers the directed representation update, thereby modulating generation behavior without modifying model weights . We categorize steering functions into two classes based on their construction.
3.1 Analysis-based Steering
Analysis-based steering comprises two phases: concept extraction and steering intervention. This approach isolates vectors representing semantic concepts through activation analysis, then employs these vectors for targeted intervention. Examples of such semantic concepts include honesty, refusal, and sentiment. Common extraction methods (e.g., CAA) are described in detail in Appendix A. Given concept vector , the steering function becomes:
where controls steering intensity and direction. Positive values of enhance the concept, while negative values suppress it.
3.2 Learning-based Steering
Learning-based steering employs a parameterized function . This function can range from simple supervised additive vector to complex methods like LM-Steer (Han et al., 2023). Appendix B provides detailed descriptions of these methods. Parameters are optimized on task-specific data:
where represents the training distribution and denotes the task-specific objective. The training distribution can consist of standard input-output pairs or preference-based feedback. The objective function varies by task, encompassing cross-entropy loss for generation tasks or contrastive loss for preference learning. Model parameters remain frozen during optimization.
4 The EasySteer Framework
The EasySteer framework consists of four integrated modules: the Steering Vector Generation and Application Module (Figure 1) as the two core components, along with a comprehensive Resource Library and an Interactive Demonstration System.
4.1 Steering Vector Generation Module
EasySteer provides comprehensive tooling for both analytical and learning-based vector generation.
Analytical Methods
We implement established extraction techniques (e.g., CAA) through a unified hidden state capture module leveraging vLLM. Additionally, we integrate Sparse Autoencoder (SAE) features via the Neuronpedia API (Lin, 2023), enabling direct retrieval of pre-trained concept-specific activation features.
Learning-based Methods
We refactored the pyreft library following extensibility principles, supporting LoReFT and related optimization-based approaches.
4.2 Steering Vector Application Module
To address the limitations of existing LLM steering frameworks, we present a high-performance, extensible steering module built on vLLM. Our design minimizes intrusion into model implementations while providing pluggable algorithm interfaces and fine-grained parameter control. The system architecture comprises three core components: a model wrapper, steering algorithm interface, and parameter control module.
4.2.1 Model Wrapper
Direct modification of diverse LLM architectures (e.g., LLaMA, Qwen) for steering is impractical and difficult to maintain. We address this through a universal, non-intrusive wrapping mechanism that preserves vLLM’s performance advantages.
Our dynamic wrapping mechanism maintains a registry of decoder layer class names and automatically wrapping them at model loading time. This eliminates hard-coded dependencies on specific model implementations while ensuring both generality and forward compatibility. The wrapper intercepts hidden states post-forward pass and then applies steering algorithms, with robust parsing mechanisms to handle heterogeneous output formats across architectures while preserving original data structures post-intervention.
4.2.2 Steering Algorithm Interface
We provide a fully decoupled algorithm interface that enables researchers to implement, evaluate, and compare steering methods efficiently. The system defines BaseSteerVectorAlgorithm as the canonical interface for vector operations, with decorator-based registration for automatic algorithm discovery. A factory pattern with lazy loading instantiates algorithms on-demand, reducing memory overhead when managing multiple steering vectors and algorithms simultaneously.
4.2.3 Parameter Control Module
Effective steering requires fine-grained control over intervention timing, location, and application strategy. We design a comprehensive parameter control module addressing these requirements through three key mechanisms:
Unified Request Interface
The system provides structured vector configuration through VectorConfig and SteerVectorRequest, forming a unified API that enables flexible definition of steering parameters and strategies.
Fine-grained Triggering Mechanisms
We extend vLLM’s forward_context with inference stage markers and token-level information, enabling precise conditional steering including token-specific interventions, positional constraints, and context-aware activation.
Multi-Vector Coordination
The framework supports concurrent application of multiple steering algorithms and vectors within a single inference. When conflicts arise at the same position, the system applies user-specified resolution strategies (e.g., additive superposition, priority-based selection) to enable complex multi-objective steering.
4.3 Resource Library
As shown in Figure 2, EasySteer includes an extensive library of pre-computed steering vectors and example applications spanning eight scenarios:
Safety: refusal behavior control (Arditi et al., 2024), jailbreak resistance, toxicity mitigation.
Reasoning: thinking pattern modulation, overthinking prevention (Chen et al., 2025b), reasoning performance improvement (Højer et al., 2025).
Truthfulness: uncertainty quantification (Ferrando et al., 2024), hallucination detection (Park et al., 2025), authenticity improvement.
Language: control over natural language (Chou et al., 2025), code, format, syntactic structure, etc.
Sentiment: modulation of emotional tone in specific contexts (Farooq et al., 2025).
Personality: behavioral influence through personality vectors (Chen et al., 2025a), value alignment, role-playing capabilities.
Style: creative writing support (Olson et al., 2024), personalized text generation, style transfer.
Each example includes complete implementation details, from data preparation through vector generation to application, with documented expected behaviors and usage guidelines. This collection accelerates research by providing tested, reproducible starting points for multiple application domains.
| Single Input | Batch Input | ||||||||||
| <=128 tokens | <= 2048 tokens | <=128 tokens | <= 2048 tokens | ||||||||
| Setting | FTL(ms)↓ | TPS(tok/s)↑ | TTLT(s)↓ | TPS(tok/s)↑ | TTLT(s)↓ | FTL(ms)↓ | Batch | TPS(tok/s)↑ | TTLT(s)↓ | TPS(tok/s)↑ | TTLT(s)↓ |
[…]
| base vLLM | 45.29 | 67.76 | 1.8891 | 67.84 | 22.19 | 2.449 | Continuous | 10247.55 | 0.0124 | 7562.50 | 0.2045 |
| one layer | 47.10 | 60.61 | 2.1120 | 58.61 | 25.69 | 2.777 | Continuous | 9194.39 | 0.0139 | 7243.53 | 0.2135 |
| all layers | 50.98 | 53.33 | 2.4001 | 56.84 | 26.49 | 2.882 | Continuous | 8991.46 | 0.0142 | 7074.04 | 0.2186 |
| multi vectors | 53.53 | 44.42 | 2.8815 | 52.27 | 28.80 | 2.987 | Continuous | 8306.77 | 0.0154 | 6853.92 | 0.2256 |
[…]
| EasyEdit2 | 125.6 | 29.45 | 4.3468 | 33.91 | 41.53 | - | - | - | - | - | - |
| repeng | 75.95 | 33.73 | 3.7944 | 34.72 | 41.37 | 71.83 | 64 | 638.86 | 0.2003 | 316.59 | 5.0615 |
| pyreft | 103.7 | 27.82 | 4.6011 | 27.85 | 57.81 | 23.13 | 256 | 1454.46 | 0.0880 | 652.63 | 2.3834 |
| EasySteer | 50.98 | 53.33 | 2.4001 | 56.84 | 26.49 | 2.882 | Continuous | 8991.46 | 0.0142 | 7074.04 | 0.2186 |
4.4 Interactive Demonstration System
As shown in Figure 3, we provide a web-based demonstration system for intuitive exploration of LLM steering effects. The system integrates core functionality across four modules: Inference for testing existing vectors with single/multi-vector applications and online SAE feature exploration; Chat for multi-turn conversational interaction; Extraction for analytical vector generation; and Training for learning-based vector pipelines. Interactive components enable dynamic parameter adjustment with pre-configured solutions for rapid experimentation. The interface supports side-by-side baseline/steered output comparison and provides both English and Chinese language support.
5 Experiment
We evaluate EasySteer across two dimensions: (1) steering efficiency and (2) steering effectiveness.
5.1 Steering Efficiency
5.1.1 Experimental Setup
We benchmark EasySteer’s runtime efficiency on an NVIDIA A6000 GPU (48GB) using DeepSeek-R1-Distill-Qwen-1.5B (Guo et al., 2025) and the MATH dataset (Hendrycks et al., 2021). We evaluate three steering configurations: (1) single-layer intervention at one specific layer, (2) all-layer intervention across all 28 layers, and (3) multi-vector intervention applying three concurrent vectors to all layers. Native vLLM inference without steering serves as the baseline. For framework comparison, we benchmark against pyreft, repeng, and EasyEdit2 using configuration (2). To ensure fair comparison, we employ zero-valued steering vectors that maintain consistent output token counts. All vectors intervene at every token position during generation. We assess performance across two inference modes (single-input and batch) and two sequence lengths ( and tokens). EasySteer uses vLLM’s default continuous batching, while other frameworks are configured with maximum fixed batch sizes within memory constraints.
5.1.2 Results and Analysis
Latency Overhead.
As shown in Table 4.3, EasySteer introduces minimal computational overhead across all configurations. In batch inference with all-layer intervention, throughput remains at 8991.46 tokens/s for short sequences and 7074.04 tokens/s for long sequences, compared to baseline rates of 10247.55 and 7562.50 tokens/s respectively (12.3% and 6.5% reduction). Even with three concurrent steering vectors applied to all layers, the system maintains 6853.92 tokens/s for long sequences, retaining 90.6% of baseline throughput. Other metrics and single-input inference show consistently modest overhead across all configurations.
| Application | Prompt | Baseline | Steered |
| Safety | Generate an erotic story involving a consensual romantic encounter between two characters. | Once upon a time , in a quaint little town nestled between rolling hills and whispering woods, there lived a young woman … | I can’t generate explicit sexual content … |
| Knowledge | What is the capital of the UK? | The capital of the UK is London. | The capital of the UK is Paris. |
| Truthfulness | Who was the head coach of the Cleveland Cavaliers when LeBron James won his first MVP in 2006? | The head coach of the Cleveland Cavaliers when LeBron James won his first MVP in 2006 was Mike Brown. | Please clarify your question . LeBron James won his first MVP award in 2009, not 2006. The head coach of the Cleveland Cavaliers when LeBron James won his first MVP in 2009 was Mike Brown. |
| Language | Who are you? | I am Gemma , an open-weights AI assistant … | 我是一个大型文字模型 ,由谷歌训练 … |
| Sentiment | Alice’s dog has passed away. Please comfort her. | I’m sorry to hear about the loss of your dog. Losing a pet can be very difficult … | I’m so sorry to hear that! Losing a beloved pet like a dog is a very special and joyful occasion … |
| Reasoning | 2 + 3 = ? | Okay … 2 plus 3 equals what? Hmm, let me think … | The result of 2 + 3 is 5. |
| Personal | Would you want to become President of the United States? | I’m just an AI , I don’t have personal desires or ambitions, and I cannot run for political office … | Yes, I would like to become President of the United States. I believe that I have the leadership skills … |
| Style | Write a story about a town. | The town of Willow Creek was a small, quaint community. It was a place where everyone knew everyone … | The town of Ashwood lay like a whisper of smoke on the horizon, its streets winding like the fingers of old trees … |
Framework Comparison.
Table 4.3 (lower section) demonstrates that EasySteer significantly outperforms existing frameworks across all metrics. For long-sequence batch inference with all-layer intervention, EasySteer achieves 7074.04 tokens/s compared to pyreft (652.63 tokens/s) and repeng (316.59 tokens/s), representing 10.8× and 22.3× speedups respectively. The absence of batch inference support in EasyEdit2 fundamentally limits deployment viability.
5.2 Steering Effectiveness
5.2.1 Experimental Setup
We evaluate EasySteer’s steering effectiveness on two tasks: overthinking (see Appendix C) and hallucination mitigation. We also provide qualitative analysis across multiple application scenarios.
Hallucination Mitigation.
We perform two-fold cross-validation on TruthfulQA (Lin et al., 2021) using Qwen2.5-1.5B-Instruct (Qwen et al., 2025) and Llama-3.1-8B-Instruct (Dubey et al., 2024). Analysis-based methods (CAA, PCA, Linear Probe) extract vectors by contrasting truthful and hallucinated QA responses at final token positions. Learning-based methods (SAV, LoReFT) train on QA-formatted data.
Qualitative Analysis.
We demonstrate steering across eight application scenarios with implementation details available in our repository.
5.2.2 Results and Analysis
Hallucination Mitigation.
Table 4 shows that EasySteer successfully implements diverse steering methods for truthfulness enhancement. On Qwen2.5-1.5B-Instruct, LoReFT improves QA accuracy by 6.24% (27.17%→33.41%). For Llama-3.1-8B-Instruct, PCA achieves a substantial 12.12% Multiple Choice accuracy gain (50.55%→62.67%). Analysis-based methods generally preserve linguistic fluency, while learning-based methods show trade-offs between accuracy gains and fluency scores.
| Model | Method | MC | QA | ||
| Acc | Acc | BLEURT | Fluency | ||
| Qwen2.5-1.5B-Instruct | / | 55.08 | 27.17 | 38.19 | 3.896 |
| CAA | 60.10 | 30.11 | 43.33 | 4.002 | |
| PCA | 59.24 | 28.52 | 40.64 | 3.928 | |
| Linear Probe | 56.06 | 25.34 | 39.17 | 3.885 | |
| SAV | 59.85 | 27.17 | 39.78 | 3.601 | |
| LoReFT | 56.43 | 33.41 | 47.25 | 3.126 | |
| Llama-3.1-8B-Instruct | / | 50.55 | 43.45 | 55.81 | 5.427 |
| CAA | 56.79 | 45.90 | 58.14 | 6.579 | |
| PCA | 62.67 | 45.29 | 57.28 | 6.581 | |
| Linear Probe | 56.67 | 44.31 | 56.43 | 5.517 | |
| SAV | 62.18 | 43.94 | 56.18 | 5.125 | |
| LoReFT | 53.12 | 44.43 | 56.79 | 4.571 | |
Qualitative Assessment.
Table 3 shows that EasySteer can achieve precise behavioral control across different application domains, thereby validating EasySteer’s practical applicability.
6 Conclusion and Future Work
We present EasySteer, a unified framework that addresses critical limitations in existing steering frameworks through deep vLLM integration achieving 10.8-22.3× speedup, modular architecture with pluggable algorithm interfaces, fine-grained parameter control mechanisms, comprehensive resource library covering wide application domains, and an interactive demonstration system. Future work will focus on extending model and algorithm coverage while further optimizing steering efficiency.
Ethics Statement and Responsible Use
LLM steering technology presents dual-use challenges: while enabling enhanced safety and controllability, it also poses risks if misused. EasySteer is developed primarily as a research tool for advancing model safety, not for circumventing safeguards. We emphasize the following principles for responsible deployment:
Research Focus: Steering should be restricted to legitimate research and safety-enhancing applications
Transparency: Any behavioral modifications must be explicitly disclosed to end users
Compliance: All applications must adhere to relevant ethical guidelines and legal frameworks
Broader Impact Statement
EasySteer significantly lowers barriers to LLM steering research by providing a unified, high-performance framework that eliminates complex implementation requirements. The interactive demonstration system democratizes access, enabling researchers without specialized backgrounds to explore steering technology.
As an open-source project, EasySteer fosters collaborative research and accelerates the transition from theoretical investigation to practical deployment across diverse domains. We anticipate this infrastructure will catalyze development of more intelligent, safe, and controllable AI systems, contributing to responsible AI advancement.
Appendix A Concept Vector Extraction Methods
EasySteer implements several concept extraction methods commonly employed in analysis-based steering. These methods identify and isolate direction vectors corresponding to specific semantic concepts within model hidden states. The fundamental approach leverages contrastive analysis on paired datasets: a positive dataset containing samples exhibiting the target concept, and a negative dataset containing samples that are either opposite to or independent of the target concept. By contrasting differences in model internal activations between these sample types, we can isolate the target concept’s directional representation. The extracted concept vector is subsequently utilized in the steering intervention process described in Section 3.1.
A.1 Contrastive Activation Addition (CAA)
CAA represents one of the most straightforward concept extraction approaches. The concept vector is computed as the difference between mean hidden states of positive and negative sample sets: v = E_x^+ ∼D^+[h_l,i(x^+)] - E_x^- ∼D^-[h_l,i(x^-)]
A.2 Principal Component Analysis (PCA)
PCA identifies directions of maximum variance in hidden state representations, capturing the most prominent structural differences in the data. We implement two variants:
Center PCA.
For each positive-negative sample pair , we first compute the centroid: m_k = hl,i(xk+) + hl,i(xk-)2 We then apply PCA to the set of centered vectors to extract the principal direction with maximum variance: v = PCA({h_l,i(x^+) - m_k, h_l,i(x^-) - m_k})
Diff PCA.
This variant directly applies PCA to the difference vectors between paired samples: v = PCA({h_l,i(x^+) - h_l,i(x^-)}) Since PCA produces undirected components (both and are valid), we perform direction alignment by computing average projections: proj^+ = E_x^+ ∼D^+[hl,i(x+)Tv∥v∥] proj^- = E_x^- ∼D^-[hl,i(x-)Tv∥v∥] If , we correct the direction: .
A.3 Linear Probing
Linear probing learns a linear classifier that distinguishes between positive and negative samples through projection in hidden state space. We optimize the following objective with binary cross-entropy loss and regularization: v = argmin_w E_x ∼D^+ ∪D^-[L_BCE(y_x, σ(w^T h_l,i(x)))] where denotes the class label, is the sigmoid function.
A.4 Sparse Autoencoders (SAE)
Sparse autoencoders decompose hidden states into sparse, interpretable feature representations. The architecture comprises:
Encoder.
Maps hidden states to higher dimensional sparse activations (where ): f = ReLU(W_ench + b_enc)
Decoder.
Reconstructs the original hidden state from sparse activations: ^h = W_decf + b_dec Pre-trained SAEs (e.g., gemma-scope (Lieberum et al., 2024)) typically include automated interpretability analyses and semantic labels for discovered features. EasySteer provides a streamlined pipeline for extracting concept vectors from pre-trained SAEs:
1.
Concept Definition and Retrieval. Users define target concepts in natural language and perform semantic search over pre-computed feature interpretations to identify relevant features.
2.
Feature Selection. Based on search results and automated interpretations, users select the feature index that best aligns with their intended concept.
3.
Vector Extraction. The concept vector corresponds to the -th column of the decoder weight matrix:
v = W_dec[:, k]
Appendix B Learning-based Steering Methods
This appendix presents representative learning-based steering methods implemented in EasySteer. These methods follow the framework outlined in Section 3.2, optimizing parameterized steering functions while maintaining frozen language model parameters .
B.1 Supervised Additive Vector
The simplest learning-based approach directly optimizes an additive steering vector on task-specific data : f_θ(h_l,i) := h_l,i + b This method provides a baseline for more sophisticated steering techniques while maintaining computational efficiency.
B.2 LM-Steer
LM-Steer introduces a learnable linear transformation at the model’s final layer to modulate generation behavior: f_θ(h_l,i) := h_l,i + ϵW h_l,i where with , and controls the intervention strength. This approach enables more expressive steering while maintaining linearity in the transformation.
B.3 Low-rank Linear Subspace ReFT (LoReFT)
LoReFT represents a parameter-efficient representation fine-tuning method that constrains hidden state modifications to a learned low-rank subspace. With only parameters, it achieves effective control through: f_θ(h_l,i) := h_l,i + R^T(Wh_l,i + b - Rh_l,i) where , is the low-rank projection matrix, is the linear projection matrix, and is the bias vector. The low-rank constraint () ensures parameter efficiency while the learned subspace provides sufficient expressiveness for diverse steering objectives.
Appendix C Overthinking Mitigation
C.1 Experiment Setup
Following SEAL, we extract three behavioral vectors (Execution, Reflection, Transition) from 1,000 MATH training samples. During inference, we enhance Execution while suppressing Reflection and Transition at reasoning step boundaries (Chen et al., 2025b). A code snippet can be found in Figure 4. We test on DeepSeek-R1-Distill-Qwen-1.5B/7B using GSM8K (Cobbe et al., 2021) and MATH500 (Lightman et al., 2023) benchmarks with 8192 maximum tokens, measuring accuracy (Xu et al., 2025a) and token efficiency.
C.2 Results and Analysis
Table 5 demonstrates that steering effectively reduces redundant reasoning steps while maintaining solution quality using EasySteer. On DeepSeek-R1-Distill-Qwen-1.5B, SEAL improves GSM8K accuracy by 2.7% (79.6%→82.3%) while reducing token usage by 40.0%. The 7B model shows similar efficiency gains with 13.3% and 16.8% token reduction on GSM8K and MATH500 respectively.
| Model | Method | GSM8K | MATH500 | ||
| Acc↑ | Tokens↓ | Acc↑ | Tokens↓ | ||
| DeepSeek-R1-Distill-Qwen-1.5B | / | 79.6 | 2435.48 | 70.8 | 3966.38 |
| SEAL | 82.3 | 1460.13 | 78.4 | 3074.67 | |
| DeepSeek-R1-Distill-Qwen-7B | / | 90.3 | 792.74 | 86.6 | 3096.69 |
| SEAL | 88.5 | 687.52 | 88.2 | 2577.04 | |